Two machines work together and complete work in 2 hours 24 minutes. If one machine works on its own it takes 2 hours longer than the other machine.How long does it take the slower machine to complete the work on its own??
Answer by stanbon(75887)

Two machines work together and complete work in 2 hours 24 minutes. If one machine works on its own it takes 2 hours longer than the other machine.How long does it take the slower machine to complete the work on its own??
-----
Together rate: 1/144 job/min
Slower mach. rate: 1/(x+120) job/min
Faster mach. rate: 1/x job/min
-----
Equation:
rate + rate = together rate:
1/x + 1/(x+120) = 1/144
------
144(x+120) + 144x = x(x+120)
---
144x + 144*120 + 144x = x^2 + 120x
------
x^2 - 168x - 144*120 = 0
-------------------------
x = 240 minutes (time for the faster worker)
x+120 = 360 min (time for the slower worker)
============================
Cheers,
Stan H.
=============